OpenStudy (anonymous):

Write the equation of the line tangent to the circle at the given point. x2 + y2 = 13 at (2, 3) Question 14 options: y=−23x−133 y=−23x+133 y=23x−133 y=23x+133

OpenStudy (anonymous):

The circle's origin at (0,0) and radius \(r=\sqrt{13}\)... a line from the origin to (2,3) is perpendicular to the line tangent to the circle at given point... and the slope from (0,0) to (2,3) is \[m=\frac{0-3}{0-2}=\frac{3}{2}\]so the line tangent to the given circle will have a slope \[m_1=- \frac{1}{m_2}\] for perpendicular lines... so if \(m_2=m=3/2\), then \(m_1=-2/3\). Since a point (2,3) is the tangent point of the tangent line, by point-slope form \[(y-y_1)=m(x-x_1)\]\[(y-3)=-\frac{2}{3}(x-2)\]\[3(y-3)=-2x+4\]\[3y-9=-2x+4\]\[3y=-2x+4+9=-2x+13\]\[y=-\frac{2}{3}x+\frac{13}{3}\] it seems the second one...